OpenStudy (fanduekisses):

Why is the cosine and sine of -5pi/3 (1/2, sqrt(3)/2 respectively?

OpenStudy (fanduekisses):

I thought it'd be: \[-\frac{ 1 }{ 2 }, \frac{ -\sqrt{3} }{ 2 }\] since 5pi/3 is in the 3rd quadrant and both cosine and sine are negative.

OpenStudy (fanduekisses):

@SolomonZelman

Nnesha (nnesha):

no look at the unit circle again

OpenStudy (fanduekisses):

Oh yeah I mean 1/2, -sqrt(3)/2

Nnesha (nnesha):

it's negative 5pi/3 but you're thinking about positive 5pi/3

Nnesha (nnesha):

the easy way is add 2pi to make it positive -5pi/3+2pi = pi/3 which is in first quadrant so that's why

OpenStudy (fanduekisses):

how does the negative sign affect it?

Nnesha (nnesha):

to*

OpenStudy (fanduekisses):

ohh like coterminal?

Nnesha (nnesha):

yeaho

OpenStudy (fanduekisses):

ohh ok thank you!

OpenStudy (fanduekisses):

what about sin t= 4/5 sin(pi - t) and sin(t + pi)

jimthompson5910 (jim_thompson5910):

The negative angle tells you to go in a clockwise direction instead of counter-clockwise like normal 

Nnesha (nnesha):

np :) good work! for me its easy to add 2pi or 360 but some people like to keep going around the circle

Nnesha (nnesha):

yea that ^^

jimthompson5910 (jim_thompson5910):

`what about sin t= 4/5` `sin(pi - t) and sin(t + pi)` use these identities sin(A+B) = sin(A)*cos(B) + cos(A)*sin(B) sin(A-B) = sin(A)*cos(B) - cos(A)*sin(B) sin(pi) = 0 cos(pi) = -1 You'll need to figure out what cos(t) is equal to. Hopefully it tells you which quadrant angle t is in?